2-48 A refinisher of antiques named Constance has beern so successful with her small business that she is plan- ning to expand her shop. She is going to start enlarg ing her shop by purchasing the following equipment. A Cost of Power New Original Original Sizing Equipment Equipment Size Equipment Exponent Size Varnish bath 50 gal $3500 0.80 75 gal Power scraper 3/4 hp $250 0.22 1.5 hp Paint booth 3 ft3 $3000 0.6 12 ft3 What would be the net cost to Constance to obtain this equipment? Assume that she can trade the old equipment in for 15% of its original cost. Assume there has been no inflation in equipment prices.
Answer #1
Equipment Cost of old equipment (C) Original size (Q1) New size(Q2)

Q=(Q2/Q1)

Power sizing exponent

(n)

(Q)n Cost of new equipment = C(Q)n
Varnish bath 3500 50 gal 75 gal 1.5 0.8 1.38 4830
Power scrapper 250 0.75 hp 1.5 hp 2 0.22 1.16 290
Paint booth 3000 3 ft3 12 ft3 4 0.6 2.3 6900

Total cost of new equipment = 4830+290+6900 = 12020$

Old equipment total cost = 6750$

Sale value of old equipment = 6750*15% = 1012.5$

Net cost to obtain new equipment = 12020-1012.5 = 11007.5$